OpenStudy (anonymous):

find the domain of function g(x)=x+9/x^3+4x. Can someone show me how to work this one step by step please?

OpenStudy (anonymous):

someone please help me with this problem

OpenStudy (anonymous):

When finding the domain of a function you just need to make sure that everything is defined. In your case there's only division by zero to worry about if you write this as \[\frac{ (x+9) }{ x(x^{2}+4) }\] you can see that the only problematic point is x=0 as there's no solution for x^2+4=0 for real numbers.

OpenStudy (anonymous):

so your final solution would be R without 0

OpenStudy (anonymous):

what do you mean R without 0

OpenStudy (anonymous):

R is the symbol for real numbers or should i say if x is from R then x can be any real number. On the other hand since we don't know how to divide by zero we need to exclude that case. Not sure if that makes any sense to you but if not you should read up on numbers on wikipedia.

OpenStudy (anonymous):

yes i see now. Thank you very much

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can you help me with another similar one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

find the domain of function g(x)= 7x/x^2-4

OpenStudy (anonymous):

can you show me how to enter the problem in ordered pair solution. exp. ( )

OpenStudy (anonymous):

i guess using the infinity sign

OpenStudy (anonymous):

and the big u symbol, which i am not sure what that means

OpenStudy (anonymous):

this time it's 7x/(x-2)*(x+2) so we excluse 2 and -2 for ordered pairs you can write it as any of the following R\{2,-2} (-\[(-\infty,-2)\cup(-2,2)\cup(2,\infty)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

the U symbol means union basically

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

for example {3,5)U{1,2,3}={1,2,3,4,5}

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

another one ?

OpenStudy (anonymous):

sure